# Robust Facial Reactions Generation: An Emotion-Aware Framework with Modality Compensation

### v1

@@ 1 Introduction
In recent years, the field of human-computer interaction has increasingly focused on developing systems capable of comprehending and appropriately responding to human conversations. This progression is crucial for enabling more natural and effective communication between humans and machines. A pivotal aspect of this domain involves enabling virtual agents and robots to autonomously generate realistic, human-like facial reactions that are both contextually appropriate and diverse in response to their human interlocutors’ multi-modal behaviours, typically encompassing speech and facial expressions [55, 63, 51, 49, 43, 13, 28, 30, 11, 26, 27, 17, 62, 9, 21, 33, 34, 14, 35, 15, 31, 25, 19].
Recent developments in Multiple Appropriate Facial Reaction Generation (MAFRG) have overcome the limitations of traditional one-to-one mapping methods, which struggled with the variability in appropriate reactions. By reframing the problem as a mapping from speaker behaviour to a distribution of appropriate reactions, these models have enhanced the diversity and synchrony of generated facial reactions [40, 37, 45, 39, 61, 44, 18, 32, 16].
Despite significant advancements in leveraging multi-modal inputs, current methodologies often fall short in two critical areas. Firstly, existing models typically assume the constant availability of all modalities—audio and visual. This assumption does not hold in real-world scenarios where one or more modalities might be intermittently unavailable due to factors such as network issues, environmental noise, or occlusions [41] [58]. Secondly, there is insufficient emphasis on the speaker’s emotional context, which is crucial for generating appropriate listener reactions [50] [38]. Effectively incorporating emotional cues into the facial reaction generation process is vital for the authenticity and relevance of the produced responses [7] [59]. The failure to address these challenges results in systems that are less effective and robust in practical applications [1, 8, 3, 10, 4, 2]. There is a clear need for a framework that can both gracefully handle missing modality information and effectively integrate emotional cues.
@@ 2 Preliminaries
Most existing MAFRG models typically encompass four stages, as illustrated in Figure 1: Unimodal Feature Encoding Stage, Multimodal Feature Fusion Stage, Decoding Stage, and Visualisation Stage [48, 42, 56, 57, 29, 8]. This structure ensures the learning of appropriate and diverse listener facial representations [53]. The Unimodal Feature Encoding Stage generally corresponds to the visual/audio unimodal encoder $E_{f}/E_{s}$, the Multimodal Feature Fusion Stage to the multimodal fusion module $E$, the Decoding Stage to the multiple appropriate facial reaction representations decoder $D$, and finally, the Visualisation Stage to a visualisation module VIS. This section delineates this standard model structure.

### v2

@@ 1 Introduction
In recent years, the field of human-computer interaction has increasingly focused on developing systems capable of comprehending and appropriately responding to human conversations. This progression is crucial for enabling more natural and effective communication between humans and machines. A pivotal aspect of this domain involves enabling virtual agents and robots to autonomously generate realistic, human-like facial reactions that are both contextually appropriate and diverse in response to their human interlocutors’ multi-modal behaviours, typically encompassing speech and facial expressions [27, 32, 23, 21, 16, 5].
Recent developments in Multiple Appropriate Facial Reaction Generation (MAFRG) have overcome the limitations of traditional one-to-one mapping methods, which struggled with the variability in appropriate reactions. By reframing the problem as a mapping from speaker behaviour to a distribution of appropriate reactions, these models have enhanced the diversity and synchrony of generated facial reactions [14, 11, 18, 13, 31, 17].
Despite significant advancements in leveraging multi-modal inputs, current methodologies often fall short in two critical areas. Firstly, existing models typically assume the constant availability of all modalities—audio and visual. This assumption does not hold in real-world scenarios where one or more modalities might be intermittently unavailable due to factors such as network issues, environmental noise, or occlusions [15] [28]. Secondly, there is insufficient emphasis on the speaker’s emotional context, which is crucial for generating appropriate listener reactions [22] [12]. Effectively incorporating emotional cues into the facial reaction generation process is vital for the authenticity and relevance of the produced responses [3] [29]. The failure to address these challenges results in systems that are less effective and robust in practical applications. There is a clear need for a framework that can both gracefully handle missing modality information and effectively integrate emotional cues.
@@ 2 Preliminaries
Most existing MAFRG models typically encompass four stages, as illustrated in Figure 1: Unimodal Feature Encoding Stage, Multimodal Feature Fusion Stage, Decoding Stage, and Visualisation Stage. This structure ensures the learning of appropriate and diverse listener facial representations [25]. The Unimodal Feature Encoding Stage generally corresponds to the visual/audio unimodal encoder $E_{f}/E_{s}$, the Multimodal Feature Fusion Stage to the multimodal fusion module $E$, the Decoding Stage to the multiple appropriate facial reaction representations decoder $D$, and finally, the Visualisation Stage to a visualisation module VIS. This section delineates this standard model structure.